OpenStudy (anonymous):

2.What is the most precise name for quadrilateral ABCD with vertices A(-2,4), B(5,6), C(12,4), and D(5,2)? A. parallelogram B. rhombus C. quadrilateral D. rectangle 3. What is the most precise term for quadrilateral ABCD with vertices A(1,2), B(2,6), C(5,6), and D(5,3)? A. square B. rhombus C. parallelogram D. kite 9. What is the measure of one angle in a regular 16-gon? A. 2,520 B. 202.5 C. 157.5 D. 78.75 10. The measure of exterior angle of a regular 5-gon is ____ the measure of each exterior angle of a regular 9-gon? A. greater than B. less than C. equal to

OpenStudy (anonymous):

2) option d:-rectangle

OpenStudy (anonymous):

9) sum of interior angles in polygon=(n-2)180 where n=no. of sides for regular polygon, any one interior angle=[(n-2)180]/n=[(16-2)180]/16= [(14)(180)]16=157.5

OpenStudy (arbershabani97):

3) its kite, you can just draw it and easily see it

OpenStudy (arbershabani97):

10) its a, greater then, because we can identify equilateral triangle with 60-60-60, and square with 90-90-90-90, and regular 5gon 108-108-108-108-108, so interior angles increase when new lines are added, and when interior angles increase, exterior angles decrease when new lines are added (or here is the other explanation since i don't remember the formula of finding exterior angles, i'll find the interior angles 5gon --> ((n-2)*180)/n=180 ((5-2)*180)/5=108 --> so the exterior angle is 360-108=252 9gon --> ((n-2)*180)/n=180 ((9-2)*180)/9=140 --> so the exterior angle is 360-140=220 252>220, so the measure of the exterior angle of 5gon is greater then 9gon

OpenStudy (ranga):

2) is a rhombus. All sides are equal. The diagonals intersect at right angles but the diagonals are not equal in length. 10) "greater than" is the correct answer. The exterior angles of all polygons add up to 360. This is true for triangles, rectangles, pentagon, hexagon, etc... The polygon does not even have to be regular but still all external angles add to 360. External angle = 360 / n.

OpenStudy (anonymous):

I already finished this I forgot I even posted these on here. Thank ya'll though :)

OpenStudy (anonymous):

hey could you maybe post screen shots of the finial exam?